OpenStudy (song_of_the_sole):

can someone help me please im not sure how to do this i will medal!! :)

OpenStudy (song_of_the_sole):

by graphing the system of constraints find the values of x and y that maximizes the objective function

OpenStudy (song_of_the_sole):

x+y<8 2x+y<10 x>0 y>0

Directrix (directrix):

Which of these is the objective function? x+y<8 2x+y<10 x>0 y>0

OpenStudy (song_of_the_sole):

i know that x>0 and y>0 set up the graph and i would have to graph x+y<8 and 2x+y<10 and thats the part im not sure how to do is graphing it

Directrix (directrix):

We can use Demos.com to graph the 4 inequalities. What I think is that you have left off the objective function from the posted problem. We need it. These are the constraints: x+y<8 2x+y<10 x>0 y>0

OpenStudy (song_of_the_sole):

oh sorry i did lol here it is maximum for N=100x+40y

Directrix (directrix):

Look at the attachment. The problem has an objective function given at the beginning of the problem.

OpenStudy (song_of_the_sole):

ok?

Directrix (directrix):

Okay. We need the boundary points to check in the equation of the objective function. Let's graph and see what shows up. desmos.com

OpenStudy (song_of_the_sole):

im not sure how to

OpenStudy (song_of_the_sole):

okay i pulled up desmos what do i do now?

OpenStudy (johnnyy8bp):

140x 10x65

OpenStudy (song_of_the_sole):

what?

Directrix (directrix):

>>okay i pulled up desmos what do i do now? Enter the constraint inequalities. Hold on.

OpenStudy (song_of_the_sole):

what???

Directrix (directrix):

I don't see a region which is colored red, green, blue, and purple. I see an intersection at (0,0), I think.

Directrix (directrix):

Let's try to solve these algebraically

OpenStudy (song_of_the_sole):

oh they have that line under the > < btw

Directrix (directrix):

>> oh they have that line under the > < btw I don't know what you are saying by that

Directrix (directrix):

I think that this is the common restraint region. Attached

Directrix (directrix):

Of (0,8) and (2,6) and (0,0) and (5,0) which of these maximizes N = 100x + 40 y ? Test (0,8) N = 100*0 + 40*8 = 0 + 320 ------------------- You test N = 100*x + 40*y for x = 2, y = 6 Post what you get, okay? @song_of_the_sole

OpenStudy (song_of_the_sole):

100(2)+40(6)

OpenStudy (song_of_the_sole):

right

Directrix (directrix):

I just realized what you mean about the less than or greater than symbols really being less than or equal to and greater than or equal to. ≥ and ≤

Directrix (directrix):

___> 100(2)+40(6) = what?

Directrix (directrix):

I'll check the next point: (0,0) N = 100*0 + 40*0 = 0 + 0 = 0 You do the value of N for (5,0)

Directrix (directrix):

@song_of_the_sole 100(2)+40(6) = what?

OpenStudy (song_of_the_sole):

440

Directrix (directrix):

(0,8) goes to 320 (2,6) goes to 440 (0,0) goes to 0 (5,0) goes to 500 **** Maximized

OpenStudy (song_of_the_sole):

so its 5,0

OpenStudy (song_of_the_sole):

??

Directrix (directrix):

>> find the values of x and y that maximizes the objective function x = 5 and y = 0

OpenStudy (song_of_the_sole):

okay i get it now thank you so much

Directrix (directrix):

You are welcome.